OpenStudy (anonymous):

Integrate t/(t^4 + 3). Help! Should I try breaking this up .. or substitution..?

OpenStudy (anonymous):

Let \(u=t^2\), then \(\dfrac{1}{2}du=dt\): \[\frac{1}{2}\int\frac{dt}{u^2+3}\] Then a trig sub should work.

OpenStudy (anonymous):

should be \(du\) in the integral...

OpenStudy (anonymous):

Try the following: \[\int\limits_{}{}\frac{ t }{ t^4+3 }dt=\frac{ \sqrt{3} }{ 6 } \int\limits_{}{}\frac{\frac{ 2}{ \sqrt{3} } t}{ \left( \frac{ t^2 }{ \sqrt{3}} \right)^2+1 }dt=\frac{ \sqrt{3} }{ 6 }\int\limits_{}{}\frac{ d \left( \frac{ t^2 }{ \sqrt{3} } \right) }{ \left( \frac{ t^2 }{ \sqrt{3} } \right)^2+1}\]Does it look familiar to you?

OpenStudy (anonymous):

Something like\[\frac{ \sqrt{3} }{ 6 }\arctan \left( \frac{ t^2 }{ \sqrt{3}} \right)+C\]perhaps?

OpenStudy (anonymous):

As you see, rearranging things you can have an immediate integral

OpenStudy (anonymous):

Ohhhh, thanks both for your help.